OpenStudy (anonymous):

the office jet printer can copy sue's dissertation in 16 min. The LaserJet printer can copy the same document in 18 min. If the two machines work together, how long would they take to copy the dissertation?

OpenStudy (anonymous):

I need some help please

OpenStudy (anonymous):

Welcome to Openstudy! You have to make an algebraic equation to solve this. Let x=amount of time it takes if both machines are working together So we're starting with the machines copying TOGETHER at a rate of 1/x of the dissertation per minute Officejet copies at 1/22 of the dissertation per minute Laserjet copies at 1/16 of the dissertation per minute (You can simplify to 1/8) So that makes our equation: \[1\div22 + 1\div8 = 1/x\] Then we eliminate the denominators by using the LCD (8 and 22's least common denomnator are themselves) \[(8+22)\div(22*8)1/x\]We simplify \[30\div176=1/x\] Cross multiply.. \[x=176\div30\] Solve and we have... \[x=5.867\] Get it? :)